1/8X + 1/4y=5
1/16x +1/2y=7
Solve system with addition method.

Answer by nabla(475)

First, multiply the first equation by -(1/2).
That gives:
(-1/16)x-(1/8)y=-5/2
(1/16)x+(1/2)y=7
Now we can add them.
(3/8)y=9/2
y=72/6=12.
Now for x:
(1/16)x+6=7
x=16
Check:
2+3=5
1+6=7.